OpenStudy (anonymous):

ANSWER IN 5 MINUTES FOR MEDAL + FAN!! ANSWER IN 5 MINUTES FOR MEDAL + FAN!! Write the equation in slope intercept form. 4x - 6y = -36

OpenStudy (studygurl14):

slope-intercept form: y= mx + b So...solve for y 4x - 6y = -36 -6y = -4x - 36 y = 4/6x + 6 \[y= \frac{2}{3}x+6\]